OpenStudy (anonymous):

The half-life of a radioactive substance is the time required for half of a sample to undergo radioactive decay, or for the quantity to fall to half its original amount. Carbon 14 has a half-life of 5,730 years. Suppose given samples of carbon 14 weigh (fraction 5\8) of a pound and (fraction 7\8 of a pound. What was the total weight of the samples 11460 years ago (show work please)

OpenStudy (anonymous):

That's two half lives. 5/8 * 2 * 2 = 20/8 = 5/2 pounds 7/8 * 2 * 2 = 28/8 = 7/2 pounds

OpenStudy (anonymous):

I knew it was two half lives because 5730+5730 = 11460

OpenStudy (anonymous):

thanks